Exercise 9-14

On July 1, 2020, Bridgeport Aggregates Ltd. purchased 6% bonds having a maturity value of $65,000 for $67,331. The bonds provide the bondholders with a 5% yield. The bonds mature four years later, on July 1, 2024, with interest receivable June 30 and December 31 of each year. Bridgeport uses the effective interest method to allocate unamortized discount or premium. The bonds are accounted for using the FV-OCI model with recycling. Bridgeport has a calendar year end. The fair value of the bonds at December 31, 2020 and 2021, was $67,296 and $66,402, respectively. Assume fair value adjustments are recorded at year end only. Immediately after collecting interest on December 31, 2021, the bonds were sold for $66,402.

A. Prepare the journal entry at the date of the bond purchase (July 1, 2020).

B. Prepare a bond amortization schedule to December 31, 2021.

Schedule of Interest Revenue and Bond Premium Amortization
Effective-Interest Method
Date Cash Received Interest Revenue Premium Amortized Amortized Cost of Bonds
July 1, 2020 $
Dec. 31, 2020 $ $ $
June 30, 2021
Dec. 31, 2021

C. Prepare the entries and year-end entries from December 31, 2020, through to the collection of interest on December 31, 2021.

Date

Account Titles and Explanation

Debit

Credit

December 31, 2020June 30, 2021December 31, 2021

(To record collection of interest)

(To record fair value adjustment)

December 31, 2020June 30, 2021December 31, 2021

(To record collection of interest)

December 31, 2020June 30, 2021December 31, 2021

(To record collection of interest)

D. Following the three-step approach, prepare the journal entries for the sale of the bond on December 31, 2021. Include the reclassification of unrealized gains and losses to net income
Answer #1

(A) Journal entry at the date of the bond purchase:

Credit Date Account Titles & Explanation Debit 1-Jul-2020 FV-OCI Investments $ 67,331 Cash (To record bond purchase) $67,331

(B) Bond amortization schedule upto December 31, 2021:

Schedule of Interest Revenue and Bond Premium Amortization Effective-Interest Method Cash Interest Premium Amortized Date Rec

Notes:

a) $65,000 x 6% x 6/12 = $1,950 (periodical cash received)

b) $67,331 x 5% (yield) x 6/12 = $1,683

c) $67,331 - $267 = $67,064

d) $67,064 x 5% x 6/12 = $1,677

e) $67,064 - $273 = $66,791

f) $66,791 x 5% x 6/12 = $1,670

g) $66,791 - $280 = $66,511

(C) Entries and year-end entries from December 31, 2020, through to the collection of interest on December 31, 2021:

Credit $ 267 $ 1,683 $ 232 Date Account Titles & Explanation Debit 31-Dec-2020 Cash $ 1,950 FV-OCI Investments Interest Incom

(D)  Journal entries for the sale of the bond on December 31, 2021 following the three-step approach:

Debit Credit $ 1,448 $ 1,448 $ 66,402 Date Account Titles & Explanation 31-Dec-2021 FV-OCI Investments [$66,402 - ($67,296 -